OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4 ( , )

pooja195 (pooja195):

Hi ! Welcome to openstudy y=2/3x <----equation 1 y=-2/3x+4 <---equation 2 We know what y equals so lets plug equation 1 into equation 2 2/3x=-2/3x+4 solve for x

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Pooja is already helping you :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So I need to slove x

OpenStudy (anonymous):

solve*

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

how do I do that?

pooja195 (pooja195):



pooja195 (pooja195):

Finish it off

OpenStudy (anonymous):

3/4*4 is 3

pooja195 (pooja195):

Good so x=3 plug that into equation 1 y=2/3(3) 2/3(3) = ?

OpenStudy (anonymous):

2/3*3=2

pooja195 (pooja195):

good so y=2 x=3

OpenStudy (anonymous):

thats it?

pooja195 (pooja195):

Yep :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

so it (2,3)

pooja195 (pooja195):

You flipped it (3,2)

OpenStudy (anonymous):

oh! Well Thank You so much!

pooja195 (pooja195):

no problem! :)